OpenStudy (anonymous):

find all solution to the equation 7sin^2 x -14sin x +2= -5

OpenStudy (amriju):

so effectively..we have 7 sin^2 x -14 sinx +7=0...thats a quad eqn with soln sinx=1....so find its gen soln...u'll get that in any standard maths book...or the net

OpenStudy (anonymous):

So wait is that the answer? Sinx=1?

OpenStudy (amriju):

find the value of x...it probabbly is (2n+1)pi/2..most prob...i forgot the values..

OpenStudy (anonymous):

Whats n?

OpenStudy (amriju):

okk...n is a variate ( i call it that)that assumes only integral values...by a variate ...i mean that by putting various integer values of n...we get various answers each of which satisfy the eqn...let us..for example take n=0..then we get x=pi/2...which satisfies the eqn..put n=1, -1..all satisfies the eqn...if i've written the gen soln correctly..